News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 3 September 1999 5.00am EST

During the last 48 hours the daily load has continued to run as expected. The mission plan has included predominately HRC-I observations with a series of measurements for high precision determination of the optical axis. Anomalies regarding "lost Fid Lights" were explained as being occulted by the HRC shutter during focus exercises. Ground systems have been modified to accommodate this feature. Initial checkout of HRC-I is continuing and de-dithered images are reported to look good.

The third attempt to command the HETG under software commands did complete successfully. Currently the HETG remains inserted, pending the resolution of telemetry voltage reading of approximately 6 output volts from the 28 volt Motor Control Electronics-B (MCE-B) while the OTG was in motion under MCE-A control. This may well prove to be nominal phenomenon; however, caution warrants a further investigation. Pending confirmation, we will retract and repeat one more insertion/retraction cycle on the HETG. This has now been rescheduled for this evening's contact with the Observatory.

The next load will be uplinked to the spacecraft tomorrow morning. This will be about four days in length and will continue with HRC measurements. The focus of the next series of observations will be on HRC-S in preparation for the Low Energy Transmission Grating deployment scheduled for Sunday evening.





Return to Status Reports